The product of twice a number and three is the same as the difference of five times the number and 3/4. Find the number.
Answer by checkley75(3666)

2x*3=5x-3/4
6x=5x-3/4
6x-5x=-3/4
x=-3/4 answer
proof
2*-3/4*3=5*-3/4-3/4
6*-3/4=-15/4-3/4
-18/4=-18/4